# TRIPS: Trilinear Point Splatting for Real-Time Radiance Field Rendering

### v1

@@ 4.2 Quality Comparison
In Tab. 1 and Fig. 7, we compare our approach to InstantNGP [MESK22], MipNeRF-360 [$*{}^{*}$22], 3D Gaussian Splatting [KKLD23] and ADOP [RFS22]. The latter two are the closest-related point-based radiance field rendering approaches.
On the Tanks&Temples dataset, our approach achieves in average the best LPIPS score with an improvement of 20% over the second best. In PSNR and SSIM the score is on par with state-of-the-art. On the MipNeRF-360 dataset, we obtain again the best LPIPS score, however the volumetric methods and Gaussian Splatting show an improved PSNR and SSIM. The difference can be inspected in Fig. 7. For example, in row 3, the TRIPS rendering provides better sharpness with more details, but the MipNeRF and Gaussian output is overall cleaner with less noise. On the difficult boat and office scenes, we can show that our rendering pipeline, is robust to extreme input conditions.
@@ 4.3.1 Point-Size Optimization
Figure 5: The initial COLMAP reconstruction lacks point on the pedestal of the statue (top left). Our approach distibutes the few present points and increases their sizes (bottom left) thus rendering them also in lower layers (middle). Thus our pipeline can avoid distracting holes (right).
@@
4.3.2 Point Position Optimization
To test the efficiency of our trilinear point position optimization compared to the (cheaper) approximate gradients from ADOP, we added random noise (of $0.01$) to the positions of all points after training, then optimize only point positions for 100 epochs. The result can be seen in Fig. 6. Our pipeline is able to reconstruct the correct rendering from that, while ADOP’s rendering barely improves.
Table 2: Number of resolution layers used (horse scene).
3 0.121 21.24 0.818 7.10ms
4 0.108 22.40 0.826 7.35ms
5 0.106 22.76 0.826 7.42ms
6 0.105 23.06 0.829 7.50ms
7 0.105 23.10 0.826 7.54ms
8 0.104 23.34 0.828 7.61ms
4.3.3 Number of Render Layers
Due to our trilinear point rendering algorithm, increasing the number of pyramid layers has almost no negative impact on render time. As seen in Tab. 2, having 8 layers improves quality, especially with PSNR. For reference, other approaches make use of 4 [RFS22] or 5 [$*{}^{*}$20] layers and describe significant performance impacts when increasing the number of layers [RFS22].
@@ 4.3.3 Number of Render Layers
none 0.112 24.85 0.720 11.1ms 0.103 22.73 0.822 7.7ms 0.130 24.82 0.756 17.9ms
SH-net 0.119 24.88 0.724 13.3ms 0.109 22.81 0.825 8.9ms 0.112 24.46 0.752 18.5ms
SH-point 0.127 24.32 0.702 27.4ms 0.104 22.89 0.829 10.2ms 0.122 25.15 0.756 27.3ms
Figure 7: Visual comparisons.
@@ 4.3.6 Networks
ADOP-net 0.124 24.74 0.713 10.7ms
ours 0.112 24.85 0.720 4.5ms
ours+SH 0.119 24.88 0.724 5.6ms
@@ 4.3.7 Time Scaling on Number of Points
As seen in Tab. 6, TRIPS is very efficient in rendering large amounts of points. Even for our largest scene with more than 70M points, the pipeline remains real-time capable with only 15ms required for rasterization.
@@
Time 2.5ms 5.9ms 6.4ms 13.1ms 15.0ms
@@ 4.4 Rendering Efficiency
In Tab. 7, we evaluate training and rendering time for all examined methods. Our method train for around 2-4h per scene on an Nvidia A100 and renders a novel view in around 11ms on an RTX4090.
@@ 4.6 Comparison to Prior Work with Number of Points
Ours 0.123 25.12 0.746 11.1ms
ADOP 0.138 22.86 0.753 14.5ms
Gaussian (dense) 0.225 24.06 0.773 11.4ms
Gaussian (sparse) 0.264 24.61 0.776 8.6ms
@@ 5 Limitations
Furthermore, even though the temporal consistency compared to previous point rendering approaches [$*{}^{*}$20, RFS22] has been drastically improved, slight flickering can still occur in areas with too many or too little points.

### v2

@@ 4.2 Quality Comparison
In Tab. 1 and Fig. 6, we compare our approach to InstantNGP [MESK22], MipNeRF-360 [BMV22], 3D Gaussian Splatting [KKLD23] and ADOP [RFS22]. The latter two are the closest-related point-based radiance field rendering approaches.
On the Tanks&Temples dataset, our approach achieves in average the best LPIPS score with an improvement of 20% over the second best. In PSNR and SSIM the score is on par with state-of-the-art. On the MipNeRF-360 dataset, we obtain again the best LPIPS score, however the volumetric methods and Gaussian Splatting show an improved PSNR and SSIM. The difference can be inspected in Fig. 6. For example, in row 3, the TRIPS rendering provides better sharpness with more details, but the MipNeRF-360 and Gaussian output is overall cleaner with less noise. On the difficult boat and office scenes, we can show that our rendering pipeline, is robust to extreme input conditions.
Individual scores per scene can be seen in the Appendix in Tabs. 14-19. Video results are showcased at https://youtu.be/Nw4A1tIcErQ.
@@ 4.3.1 Point-Size Optimization
Figure 5: The initial COLMAP reconstruction lacks points on the pedestal of the statue (top left). Our approach distributes the few present points and increases their sizes (bottom left) thus rendering them also in lower layers (middle). Thus our pipeline can avoid distracting holes (right).
Figure 6: Visual comparisons.
@@
none 0.225 24.85 0.720 11.1ms 0.202 22.73 0.822 7.7ms 0.219 24.82 0.756 17.9ms
SH-net 0.225 24.88 0.724 13.3ms 0.203 22.81 0.825 8.9ms 0.222 24.46 0.752 18.5ms
SH-point 0.236 24.32 0.702 27.4ms 0.200 22.89 0.829 10.2ms 0.213 25.15 0.756 27.3ms
4.3.2 Point Position Optimization
To test the efficiency of our trilinear point position optimization compared to the (cheaper) approximate gradients from ADOP, we added random noise (of $0.01$) to the positions of all points after training, then optimize only point positions for 100 epochs. The result can be seen in Fig. 7. Our pipeline is able to reconstruct the correct rendering, while ADOP’s result barely improves.
Table 3: Number of resolution layers used (horse scene).
3 0.216 21.24 0.818 7.10ms
4 0.201 22.40 0.826 7.35ms
5 0.197 22.76 0.826 7.42ms
6 0.196 23.06 0.829 7.50ms
7 0.195 23.10 0.826 7.54ms
8 0.192 23.34 0.828 7.61ms
4.3.3 Number of Render Layers
Due to our trilinear point rendering algorithm, increasing the number of pyramid layers has almost no negative impact on render time. As seen in Tab. 3, having 8 layers improves quality, especially with PSNR. For reference, other approaches make use of 4 [RFS22] or 5 [ASK20] layers and describe significant performance impacts when increasing the number of layers [RFS22].
@@ 4.3.6 Networks
ADOP-net 0.236 24.74 0.713 10.7ms
ours 0.225 24.85 0.720 4.5ms
ours+SH 0.225 24.88 0.724 5.6ms
@@ 4.3.7 Time Scaling on Number of Points
As seen in Tab. 6, TRIPS is very efficient in rendering large amounts of points.
@@
Time 2.5ms 5.9ms 6.2ms 13.1ms 15.0ms
Even for our largest scene with more than 70M points, the pipeline remains real-time capable with only 15ms required for rasterization.
@@ 4.4 Rendering Efficiency
In Tab. 7, we evaluate training and rendering time for all examined methods. Our method trains for around 2-4h per scene on an Nvidia A100 and renders a novel view in around 11ms on an RTX4090. A finer breakdown of the steps involved can be found in Tab. 8.
@@
Table 8: Breakdown of the frame time for the playground scene. Our method’s "Rasterize" consists of: counting and memory allocation with 1.9ms, splatting with 2.6ms and combined sorting and blending with 1.7ms.
Method #Points Rasterize Network Tonemap In Total
ADOP 12M 3.1ms 11.0ms 0.4ms 14.5ms
Gauss. Spl. 2M 8.6ms 8.6ms
Gauss. Spl. 8M 11.4ms 11.4ms
Ours 12M 6.2ms 4.5ms 0.4ms 11.1ms
@@ 4.6 Comparison to Prior Work with Number of Points
Furthermore, our approach performs more efficiently in scenarios with large point clouds. In the dense setup, TRIPS outperforms Gaussian Splatting, as the resolution-dependant computation cost of our neural network (4.5ms at $1920\times 1080$) catches up with our more efficient point rasterizer (see Tab. 8).
@@
Ours 0.229 25.12 0.746 11.1ms
ADOP 0.233 24.86 0.753 14.5ms
Gaussian (dense) 0.283 24.06 0.773 11.4ms
Gaussian (sparse) 0.322 24.61 0.776 8.6ms
@@ 5 Limitations
Furthermore, even though the temporal consistency compared to previous point rendering approaches [ASK20, RFS22] has been drastically improved, slight flickering can still occur in areas with too many or too little points.
Our trilinear point splatting splits up points into distinct layers and as such looses depth information. Theoretically, during recombination this could create holes in solid geometry. In practice, we could not find instances of this happening except in extreme zoom-ins far outside the training data. We believe that the per-point descriptors, the point inclusion in coarse layers, and the network-based recombination are capable to combat this issue, as reflected in the rendering quality.
@@ Acknowledgements
We thank Matthias Innmann, Stefan Romberg, Michael Gerstmayr and Tim Habigt for the fruitful discussions as well as NavVis GmbH for providing the Office dataset.
@@
A Individual Tabs: MipNeRF-360 (MipNeRF-360 resolutions)
Table 10: LPIPSVGG scores for Mip-NeRF360 scenes. $\dagger$ copied from original paper [BMV22]. $\ddagger$ copied from 3D GS [KKLD23]. Image resolutions as in MipNerf-360: half resolution for indoor, quarter resolution for outdoor. Average ours: 0.176
bicycle flowers garden stump treehill room counter kitchen bonsai
[…]
Mip-NeRF 360 0.301 0.344 0.170 0.261 0.339 0.211 0.204 0.127 0.176
Mip-NeRF 360$\ddagger$ 0.305 0.346 0.171 0.265 0.347 0.213 0.207 0.128 0.179
Gaussian Spl.$\ddagger$ 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205
TRIPS(ours) 0.194 0.297 0.159 0.268 0.266 0.147 0.158 0.127 0.111
Table 11: Normalized LPIPSVGG scores: based on the original paper [ZIE18], images should be normalized between -1 and 1 (as is in every table except Appendix Tab. 10). Scored of ours with this normalization. Average ours: 0.213
bicycle flowers garden stump treehill room counter kitchen bonsai
TRIPS(ours) 0.223 0.318 0.183 0.309 0.308 0.197 0.206 0.154 0.153
Table 12: PSNR scores for Mip-NeRF360 scenes. $\dagger$ copied from original paper [BMV22]. $\ddagger$ copied from Kerbl and Kopanas et al. [KKLD23]. Image resolutions as in MipNerf-360: half resolution for indoor, quarter resolution for outdoor. Average ours: 25.94
bicycle flowers garden stump treehill room counter kitchen bonsai
[…]
Mip-NeRF 360 24.37 21.73 26.98 26.40 22.87 31.63 29.55 32.23 33.46
Mip-NeRF 360$\ddagger$ 24.305 21.649 26.875 26.175 22.929 31.467 29.447 31.989 33.397
Gaussian Spl.$\ddagger$ 25.246 21.520 27.410 26.550 22.490 30.632 28.700 30.317 31.980
TRIPS(ours) 23.466 19.439 25.384 24.174 22.044 29.066 27.002 27.662 28.710
Table 13: SSIM scores for Mip-NeRF360 scenes. $\dagger$ copied from original paper [BMV22]. $\ddagger$ copied from Kerbl and Kopanas et al. [KKLD23]. Image resolutions as in MipNerf-360: half resolution for indoor, quarter resolution for outdoor. Average ours: 0.778
bicycle flowers garden stump treehill room counter kitchen bonsai
[…]
Mip-NeRF 360 0.685 0.583 0.813 0.744 0.632 0.913 0.894 0.920 0.941
Mip-NeRF 360$\ddagger$ 0.685 0.584 0.809 0.745 0.631 0.910 0.892 0.917 0.938
Gaussian Spl.$\ddagger$ 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938
TRIPS(ours) 0.704 0.502 0.773 0.681 0.591 0.883 0.845 0.850 0.899
B Individual Tabs: Tanks and Temples
Table 14: LPIPSVGG scores for Tanks and Temples scenes (intermediate set).
playground lighthouse francis m60 train panther family horse
InstantNGP 0.581 0.477 0.472 0.414 0.527 0.410 0.456 0.437
Mip-NeRF 360 0.350 0.346 0.343 0.313 0.486 0.285 0.277 0.244
Gaussian Spl. 0.322 0.296 0.345 0.273 0.344 0.267 0.262 0.244
ADOP 0.233 0.210 0.241 0.226 0.239 0.232 0.225 0.216
TRIPS(ours) 0.229 0.208 0.221 0.208 0.223 0.207 0.202 0.194
Table 15: PSNR scores for Tanks and Temples scenes (intermediate set).
playground lighthouse francis m60 train panther family horse
InstantNGP 18.224 20.783 23.148 24.115 18.753 26.312 21.453 19.719
Mip-NeRF 360 25.200 22.379 28.266 24.743 18.674 27.428 25.326 25.659
Gaussian Spl. 24.611 21.592 25.993 26.972 20.990 27.823 24.491 23.880
ADOP 24.856 23.057 22.036 24.707 22.335 25.666 24.013 23.261
TRIPS(ours) 25.116 23.382 24.818 25.832 22.974 26.841 23.532 23.174
Table 16: SSIM scores for Tanks and Temples scenes (intermediate set).
playground lighthouse francis m60 train panther family horse
InstantNGP 0.493 0.713 0.764 0.766 0.596 0.808 0.681 0.721
Mip-NeRF 360 0.741 0.771 0.847 0.837 0.619 0.863 0.815 0.858
Gaussian Spl. 0.766 0.790 0.847 0.868 0.734 0.880 0.820 0.853
ADOP 0.753 0.796 0.827 0.843 0.755 0.844 0.775 0.817
TRIPS(ours) 0.746 0.787 0.848 0.849 0.764 0.851 0.791 0.825
C Individual Tabs: MipNeRF-360 (our resolutions)
Table 17: LPIPSVGG scores for MipNeRF-360 scenes with our resolutions (half indoor and outdoor).
bicycle flowers garden stump treehill room counter kitchen bonsai
InstantNGP 0.600 0.587 0.516 0.591 0.606 0.354 0.393 0.286 0.266
Mip-NeRF 360 0.423 0.458 0.310 0.385 0.460 0.223 0.238 0.162 0.169
Gaussian Spl 0.363 0.448 0.245 0.359 0.460 0.234 0.231 0.158 0.215
ADOP 0.319 0.409 0.259 0.376 0.422 0.241 0.264 0.221 0.223
TRIPS(ours) 0.284 0.383 0.219 0.327 0.358 0.197 0.206 0.154 0.153
Table 18: PSNR scores for MipNeRF-360 scenes with our resolutions (half indoor and outdoor).
bicycle flowers garden stump treehill room counter kitchen bonsai
InstantNGP 21.479 19.880 23.556 22.791 21.828 29.347 26.618 28.528 30.904
Mip-NeRF 360 23.541 21.082 25.887 26.219 22.525 31.711 29.425 31.351 33.222
Gaussian Spl. 24.286 20.732 25.690 26.123 22.274 30.423 28.987 30.446 27.225
ADOP 21.910 19.432 23.711 23.700 20.312 25.975 23.088 23.614 24.330
TRIPS(ours) 22.961 19.668 25.385 24.964 21.725 29.066 27.002 27.662 28.710
Table 19: SSIM scores for MipNeRF-360 scenes with our resolutions (half indoor and outdoor).
bicycle flowers garden stump treehill room counter kitchen bonsai
InstantNGP 0.486 0.422 0.545 0.568 0.524 0.853 0.789 0.811 0.895
Mip-NeRF 360 0.635 0.522 0.730 0.727 0.611 0.909 0.882 0.901 0.940
Gaussian Spl. 0.693 0.530 0.764 0.748 0.600 0.896 0.892 0.899 0.853
ADOP 0.610 0.475 0.674 0.652 0.546 0.839 0.769 0.737 0.818
TRIPS(ours) 0.668 0.482 0.751 0.707 0.587 0.883 0.845 0.850 0.899